OpenStudy (anonymous):

Find f(x)·g(x) if f(x) = 3x and g(x) = x + 1

OpenStudy (cgreenwade2000):

Its just multiplying the two functions

OpenStudy (anonymous):

so 3x*x+1 ?

OpenStudy (cgreenwade2000):

Yes, but make sure you distribute (3x)(x+1) makes it a little clearer to see

OpenStudy (cgreenwade2000):

What you did to get 4x+1 was addition. You want to use multiplication here.

OpenStudy (cgreenwade2000):

You can rewrite it as 3x * x + 3x* 1

OpenStudy (cgreenwade2000):

Make sure you multiply 3x by both terms in g(x). That's distributing. For example 2x2=4, we know that. Well, I could also write it as 2*(1+1) If I don't distribute here, I'd get 2+1 which is 3. That's not right. So instead multiply 2 by both 1's and get 2*2=2*(1+1)=2+2=4

OpenStudy (cgreenwade2000):

Yeah, you got it, 3x^2 + 3x Hope you get your reply button working again